16. A driver was instructed to follow the given directions: 7.5 km, 66° north of east and 6.0 km, 25° west of north. Unfortunately, the driver traveled in the opposite direction in the second part of his trip. What is the driver’s resultant displacement?

1
Expert's answer
2021-11-23T12:30:57-0500

"x=7.5\cdot\cos66°+6\cdot\sin25°=5.586\ (km)"


"y=7.5\cdot\sin66°+6\cdot\cos25°=1.414\ (km)"


"d=\sqrt{5.586^2+1.414^2}=5.76\ (km)" . Answer